# Kaidi Zhang1, Yunmiao Zhang2, Honghai Yu2, Yan Hu1, Benyou Wang122footnotemark: 2

### v4

@@ Abstract
The study of social emergence has long been a central focus in social science. Traditional modeling approaches, such as rule-based Agent-Based Models (ABMs), struggle to capture the diversity and complexity of human behavior, particularly the irrational factors emphasized in behavioral economics. Recently, large language model (LLM) agents have gained traction as simulation tools for modeling human behavior in social science and role-playing applications. Studies suggest that LLMs can account for cognitive biases, emotional fluctuations, and other non-rational influences, enabling more realistic simulations of socio-economic dynamics. In this work, we introduce TwinMarket, a novel multi-agent framework that leverages LLMs to simulate socio-economic systems. Specifically, we examine how individual behaviors, through interactions and feedback mechanisms, give rise to collective dynamics and emergent phenomena. Through experiments in a simulated stock market environment, we demonstrate how individual actions can trigger group behaviors, leading to emergent outcomes such as financial bubbles and recessions. Our approach provides valuable insights into the complex interplay between individual decision-making and collective socio-economic patterns.
@@ 1 Introduction
Recent advances in large language models (LLMs) have showcased their remarkable capacity to capture complex patterns in language, decision-making, and social interactions [1, 2]. These capabilities have positioned LLMs as powerful instruments for agent-based simulations in the study of human behavior, supporting a wide range of social science applications—including conversational dynamics [3, 4, 5], emotional responses [6], and creative reasoning [7, 8]. These developments open new frontiers for social science research, enabling the modeling of fine-grained individual behavior at the micro-level and the simulation of complex, emergent phenomena at the macro-level.
@@
Our research makes three key contributions: (1) Real-World Alignment: Our framework is grounded in established behavioral theories and calibrated with real-world data, ensuring theoretical rigor and realistic human behavior modeling. This enables us to accurately reproduce certain social phenomena, allowing us to further study the mechanisms behind these phenomena. (2) Dynamic Interaction Modeling: By leveraging an LLM-based framework, we capture diverse human behaviors and their interactions, particularly in the context of information propagation. This allows us to model how opinions spread, influence decision-making, and dynamically shape collective market behaviors in real time. (3) Scalable Market Simulations: We conduct large-scale simulations to analyze the impact of group size on market behavior, providing new insights into price fluctuations and emergent dynamics.
@@ 3.1 Overall Design of TwinMarket
The design philosophy of TwinMarket emphasizes both micro-level and macro-level simulation. At the micro-level, we focus on modeling user behavior, as illustrated in Figure 2(a). This entire process is driven by BDI framework, which we will discuss in Section 3.2. At the macro-level, as illustrated in Figure 2(b), we simulate two core infrastructures of real-world financial ecosystem: an order-driven trading system that settles transactions and updates the stock index accordingly (detailed in Appendix G.4), and a social media environment that personalizes information delivery based on user behavior while enabling social interaction among agents.
@@
Table 2: Summary of data sources for TwinMarket. Further details about our data sources are provided in Appendix E.
@@
1 https://xueqiu.com/ 2 https://guba.eastmoney.com/ 3 https://data.csmar.com/
4 We choose the 50 largest and most liquid stocks on the Shanghai Stock Exchange (known as the SSE 50 index).
5 https://www.sina.com.cn/ 6 https://www.10jqka.com.cn/ 7 http://www.cninfo.com.cn/new/index.jsp
@@ 3.2 Micro-Level Simulation: BDI-driven User Modeling
In user modeling, the BDI framework plays a central role in simulating the cognitive processes of users. It enables each agent to perceive the market environment, integrate that information with its individual profile, and form a unique chain of reasoning that guides decision-making. Figure 2(a) shows how an LLM-based agent operates in a daily loop: receiving news and posts, querying from knowledge base, making trading decisions, and interacting via social platforms. Each action in this loop is governed by the agent’s beliefs, desires, and intentions, enabling nuanced and individualized behavioral responses.
@@
After completing all actions for the day, each agent evaluates its behavior based on both its actions and the environmental feedback it received. Through this process, the agent dynamically updates its beliefs, reflecting an evolving understanding and interpretation of the market. This captures the notion that a user’s perception and judgment of the environment are not static but shaped continuously by experience and new information. To track belief dynamics and explore micro-to-macro linkages, agents output a sentiment score [^2] The sentiment score ranges from 1 to 5 and is computed as the average of five self-assessed dimensions, based on user’s views on: economic fundamentals, market valuation levels, short-term market trends, the sentiment of surrounding investors, and self-confidence, as detailed in Appendix G.1. at each update step.
@@ 3.3 Macro-Level Simulation: Dynamic Social Network
Constructing Dynamic Social Network. We model the social network as a dynamic graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of users and $\mathcal{E}$ denotes weighted edges that capture similarity in trading behavior. Each user trades industry-level stock index $I=\{i_{1},\dots,i_{m}\}$, and we assume that users with more similar trading patterns are more likely to interact. To reflect temporal dynamics, we apply exponential time decay to each trading record $T(u,i,t)$, resulting in a time-weighted trading intensity $w(u,i)=\sum_{t\in T(u,i)}e^{-\lambda\cdot\Delta t}$, where $\Delta t$ is the time elapsed since trade time $t$. We then compute the weighted Jaccard similarity [^3] We use a generalized Jaccard similarity for real-valued vectors, calculated by: $S=\sum_{i}\min(w_{1},w_{2})/\sum_{i}\max(w_{1},w_{2})$. between users to determine the strength of social connections  [29]. We provide a sensitivity analysis of graph construction parameters in Appendix D.1. Figure 3 shows the dynamic social network as a chord diagram. Colors indicate user groups with similar industry preferences (e.g., Manufacturing, Finance Service), and chord thickness reflects the strength of trading similarity between groups.
@@
Information Aggregation Mechanism. To control each agent’s perception field and ensure exposure to relevant information, we design an aggregation mechanism based on the dynamic social network. For a target user $u_{t}\in\mathcal{V}$, we consider neighbors $u\in\mathcal{N}(u_{t})$ with similarity $S(u_{t},u)>\tau$. For each such user, we retrieve their posts $P(u)$ within a time window $[t_{s},t_{e}]$, where each post $p\in P(u)$ contains metadata such as timestamp $t_{p}$, upvotes $u_{p}$, and downvotes $d_{p}$. We compute a hot score $h(p)$ for each post to balance popularity and recency:
Figure 4: Hot score decay: Post influence typically lasts 2 weeks.
@@
where $t_{c}$ is the current time, $\epsilon$ prevents division by zero, and $\alpha$ is a time normalization constant controlling decay rate. As illustrated in Figure 4, we set $\alpha$ such that a post’s influence typically lasts around two weeks before its score significantly diminishes. Posts from all qualified neighbors are ranked by their hot scores, and each user is recommended the top $k$ posts per day. This mechanism enables personalized and dynamic information exposure shaped by social similarity and engagement dynamics. More technical details about information on social media are further discussed in Appendix G.2.
@@ 4.1 Micro-level Validation
We validate the micro-level realism of TwinMarket by examining emergent behavioral and economic patterns, with qualitative examples discussed in Appendix D.5.
@@ 4.2 Macro-level Validation
Figure 6 shows TwinMarket’s price change distribution (Log Return on its x-axis) mirroring real-world markets by exhibiting fat tails. While most daily price changes are small (forming its sharp central peak), large, unexpected surges or plunges (the heavy tails) occur more frequently than a normal distribution would predict, capturing actual market risk. This realism stems from our autonomous agents, which, based on BDI cognitive frameworks, process information heterogeneously, leading to diverse and potentially conflicting sentiments. These divergent sentiments then propagate through the integrated social network, becoming amplified and polarized, which in turn directly leads to extreme market outcomes.
Figure 6: Fat-tailed price changes: more extreme movements.
@@
Figure 7 shows TwinMarket capturing the leverage effect, where volatility responds asymmetrically. Specifically, initial price drops (x-axis Negative Returns) provoke larger subsequent drops (y-axis Lagged Negative Returns), thereby amplifying market volatility. This realistic asymmetry is a direct consequence of our LLM agents’ design; their BDI cognitive architectures are intentionally embedded with behavioral biases, such as loss aversion. This design makes them, akin to humans, inherently more sensitive to adverse signals. The integrated social media platform further amplifies the spread of negative sentiment, resulting in more widespread, synchronized, and potentially destabilizing collective trading behavior.
Figure 7: Initial price drops can trigger subsequent price drops.
@@
Figure 8 shows TwinMarket’s ability to replicate the volume-return relationship. Specifically, it shows that substantial increases in trading activity (x-axis Volume Change) are typically accompanied by significant market price fluctuations (y-axis Log Return). This co-movement is a key indicator, as in real markets, of collective behaviors such as herding among participants. In TwinMarket, the emergence of this co-movement provides compelling evidence demonstrates that our LLM agents, through their interactions, successfully generate collective behavioral patterns consistent with those observed in actual financial market dynamics.
Figure 8: Corr. between trading volume & price changes.
@@
Figure 9 shows TwinMarket successfully replicating volatility clustering: periods of large price swings (high Volatility on y-axis) tend to cluster together, as do periods of small swings (low Volatility). Both simulated and real volatility series clearly show these alternating groupings of high and low market turbulence over time. This emergent persistence arises as impactful market-moving events, processed via the BDI cognitive architecture, induce sustained heightened belief sensitivity. The social network then perpetuates these collective cognitive states by circulating related discussions or sentiments, prolonging periods of either erratic or stable trading.
Figure 9: Big price swings group together; small ones too.
@@
Real Data 7.26 0.14 $p{<}0.01$ 0.95
TwinMarket 5.24 0.11 $p{<}0.01$ 0.89
ABM-HPM 4.47 0.05 $p{<}0.01$ 0.82
ABM-BH 4.99 0.19 $p{<}0.01$ 0.72
We compare TwinMarket against two transitional agent-based models: the HPM Model [37], which captures investor strategy adaptation through herding behavior, predispositions, and reactions to price misalignments; and the BH Model [38, 18], which incorporates hyperbolic discounting and stochastic volatility to reflect heterogeneous investor behavior. As summarized in Table 4, TwinMarket consistently generates outputs that align more closely with real-world market statistics across all four stylized facts. This demonstrates its superior ability to replicate fundamental financial dynamics, thereby validating the empirical robustness of our simulation framework. These results were further validated for reproducibility and stability, as detailed in Appendix D.3.
The four stylized facts (SF) in Table 4 are quantified as follows: (1) Kurtosis (SF I): Measures tail heaviness; values greater than 3 indicate fat-tailed return distributions. (2) Negative Return Autocorrelation (SF II): Captures asymmetric responses to negative shocks. (3) Volume Return Correlation (SF III): A significant $p$-value ($<0.01$) suggests stronger volume–volatility linkage. (4) GARCH Volatility Clustering (SF IV): The sum $\alpha+\beta\approx 1$ reflects persistent volatility, a hallmark of financial time series [39].
@@ 5.2 Macro-level Emergence: Information Propagation
Previous work highlights how information propagation produces market-wide echo effects [42, 43]. To probe the same mechanism, we identify the day’s most salient news items and deliver them directly to users with the highest network centrality, reproducing the information asymmetry characteristic of real markets. We then rerun the experiment with a parallel set of exaggerated, rumor-like headlines (We list some examples in Appendix G.3.) to stress-test the system. Observing how these targeted, micro-level interventions cascade through the network and reshape aggregate prices and volumes allows us to treat information diffusion as a source of macro-level emergence: the collective dynamics that arise cannot be inferred by examining any single agent in isolation, but only by tracking the system-wide response to the injected signals.
Intimation & Polarization. In TwinMarket, graph edges represent behavioral similarity, meaning users closely connected exhibit stronger intimation effects. As shown in Figure 11, high-degree users receive more upvotes, indicating their greater influence within the network. These opinion leaders [44] shape decision-making, fostering homogeneity and leading to polarization, as depicted in Figure 12 (further explored in Appendix D.4). The network’s adjacency matrix highlights clusters of trading similarity, further intensified under rumor conditions. This suggests that biased or extreme information spreads more rapidly in tightly connected groups, reinforcing polarization [45, 46].
Figure 12: Heatmap of the user trading similarity adjacency matrix. Higher density indicates greater behavioral similarity.
@@ 6 Ablation Study
Table 5: Ablation study on the effects of the BDI framework and agent heterogeneity. $\uparrow$: higher is better, $\downarrow$: lower is better.
Model Variant RMSE $\downarrow$ MAE $\downarrow$ Corr. $\uparrow$ Kurtosis $\uparrow$ Lev. Effect $\uparrow$ GARCH $\alpha$uparrow GARCH $\beta$uparrow
@@
The BDI ablation notably weakens the model’s ability to generate stylized facts, such as fat tails and asymmetric volatility, highlighting the importance of cognitive reasoning in agent behavior. In contrast, removing heterogeneity leads to a sharper collapse in realism, with near-Gaussian returns and negative correlation, underscoring the necessity of strategic diversity for emergent market complexity. Together, these results confirm that both BDI framework and agent heterogeneity are essential for replicating real-world market dynamics. Additional ablation studies appear in Appendix D.2.

### v5

# Yuzhe Yang*1   Yifei Zhang*1,2   Minghao Wu*1,2 Kaidi Zhang1   Yunmiao Zhang2   Honghai Yu2,3   Yan Hu1   Benyou Wang1

@@ Abstract
Abstract: The study of social emergence has long been central to social science. Traditional modeling approaches like rule-based Agent-Based Models (ABMs) struggle to capture the diversity and complexity of human behavior, particularly the irrational factors emphasized in behavioral economics. Recently, large language model (LLM) agents have gained traction as simulation tools for modeling human behavior. Studies suggest that LLMs can account for cognitive biases, emotional fluctuations, and other non-rational influences, enabling more realistic simulations of socio-economic dynamics. In this work, we introduce TwinMarket, a novel multi-agent framework that leverages LLMs to simulate socio-economic systems. We examine how individual behaviors, through interactions and feedback mechanisms, give rise to collective dynamics and emergent phenomena. Through experiments in a simulated stock market environment, we demonstrate how individual actions trigger group behaviors, leading to emergent outcomes such as financial bubbles and recessions. Our approach provides valuable insights into the complex interplay between individual decision-making and collective socio-economic patterns.
@@ 1 Introduction
Recent advances in large language models (LLMs) have showcased their remarkable capacity to capture complex patterns in language, decision-making, and social interactions [1, 2]. These capabilities have positioned LLMs as powerful instruments for agent-based simulations in the study of human behavior, supporting a wide range of social science applications, including conversational dynamics [3, 4, 5, 6], emotional responses [7], and creative reasoning [8, 9, 10, 11]. These developments open new frontiers for social science research, enabling the modeling of fine-grained individual behavior at the micro-level and the simulation of complex, emergent phenomena at the macro-level.
@@
Our research makes three key contributions:
  • Real-World Alignment: Our framework is grounded in established behavioral theories and calibrated with real-world data, ensuring theoretical rigor and realistic human behavior modeling. This enables us to accurately reproduce certain social phenomena, allowing us to further study the mechanisms behind these phenomena.
  • Dynamic Interaction Modeling: By leveraging an LLM-based framework, we capture diverse human behaviors and their interactions, particularly in the context of information propagation. This allows us to model how opinions spread, influence decision-making, and dynamically shape collective market behaviors in real time.
  • Scalable Market Simulations: We conduct large-scale simulations to analyze the impact of group size on market behavior, providing new insights into price fluctuations and emergent dynamics.
@@ 3.1 Overall Design of TwinMarket
The design philosophy of TwinMarket emphasizes both micro-level and macro-level simulation. At the micro-level, we focus on modeling user behavior, as illustrated in Figure 2(a). This entire process is driven by BDI framework, which we will discuss in Section 3.2. At the macro-level, as illustrated in Figure 2(b), we simulate two core infrastructures of real-world financial ecosystem: an order-driven trading system that settles transactions and updates the stock index accordingly (detailed in Appendix E.5), and a social media environment that personalizes information delivery based on user behavior while enabling social interaction among agents.
@@
Table 2: Summary of data sources for TwinMarket. Further details about our data sources are provided in Appendix C.
@@ 3.2 Micro-Level Simulation: BDI-driven User Modeling
In user modeling, the BDI framework plays a central role in simulating the cognitive processes of users. It enables each agent to perceive the market environment, integrate that information with its individual profile, and form a unique chain of reasoning that guides decision-making. Figure 2(a) shows how an LLM-based agent operates in a daily loop: receiving news and posts, querying from knowledge base, making trading decisions, and interacting via social platforms. Each action in this loop is governed by the agent’s beliefs, desires, and intentions, enabling nuanced and individualized behavioral responses. A formal definition of the BDI components and the agent’s daily operational cycle is provided in Appendix E.2.
@@
After completing all actions for the day, each agent evaluates its behavior based on both its actions and the environmental feedback it received. Through this process, the agent dynamically updates its beliefs, reflecting an evolving understanding and interpretation of the market. This captures the notion that a user’s perception and judgment of the environment are not static but shaped continuously by experience and new information. To track belief dynamics and explore micro-to-macro linkages, agents output a sentiment score [^2] The sentiment score ranges from 1 to 5 and is computed as the average of five self-assessed dimensions, based on user’s views on: economic fundamentals, market valuation levels, short-term market trends, the sentiment of surrounding investors, and self-confidence, as detailed in Appendix E.1. at each update step.
@@ 3.3 Macro-Level Simulation: Dynamic Social Network
Constructing Dynamic Social Network. We model the social network as a dynamic graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of users and $\mathcal{E}$ denotes weighted edges that capture similarity in trading behavior. Each user trades industry-level stock index $I=\{i_{1},\dots,i_{m}\}$, and we assume that users with more similar trading patterns are more likely to interact [32, 33]. To reflect temporal dynamics, we apply exponential time decay to each trading record $T(u,i,t)$, resulting in a time-weighted trading intensity $w(u,i)=\sum_{t\in T(u,i)}e^{-\lambda\cdot\Delta t}$, where $\Delta t$ is the time elapsed since trade time $t$. We then compute the weighted Jaccard similarity [^3] We use a generalized Jaccard similarity for real-valued vectors, calculated by: $S=\sum_{i}\min(w_{1},w_{2})/\sum_{i}\max(w_{1},w_{2})$. between users to determine the strength of social connections [34]. We provide a sensitivity analysis of graph construction parameters in Appendix B.1. Figure 3 shows the dynamic social network as a chord diagram. Colors indicate user groups with similar industry preferences (e.g., Manufacturing, Finance Service), and chord thickness reflects the strength of trading similarity between groups.
Figure 4: Information propagation in network.
Information Aggregation Mechanism. To control each agent’s perception field and ensure exposure to relevant information, we design an aggregation mechanism based on the dynamic social network. As illustrated in Figure 4, information forms propagation chains as it spreads through the network, and the information each agent receives is restricted by its perception field. For a target user $u_{t}\in\mathcal{V}$, we consider neighbors $u\in\mathcal{N}(u_{t})$ with similarity $S(u_{t},u)>\tau$. For each such user, we retrieve their posts $P(u)$ within a time window $t_{s},t_{e}]$, where each post $p\in P(u)$ contains metadata such as timestamp $t_{p}$, upvotes $u_{p}$, and downvotes $d_{p}$. We compute a hot score $h(p)$ for each post to balance popularity and recency:
@@
where $t_{c}$ is the current time, $\epsilon$ prevents division by zero, and $\alpha$ is a time normalization constant controlling decay rate. We set $\alpha$ such that a post’s influence typically lasts around two weeks before its score significantly diminishes. Posts from all qualified neighbors are ranked by their hot scores, and each user is recommended the top $k$ posts per day. This mechanism enables personalized and dynamic information exposure shaped by social similarity and engagement dynamics. More technical details about information on social media are further discussed in Appendix E.3.
@@ 4.1 Micro-level Validation
We validate the micro-level realism of TwinMarket by examining emergent behavioral and economic patterns, with qualitative examples discussed in Appendix B.5.
@@ 4.2 Macro-level Validation
Figure 4.2 shows TwinMarket’s price change distribution (Log Return on its x-axis) mirroring real-world markets by exhibiting fat tails. While most daily price changes are small (forming its sharp central peak), large, unexpected surges or plunges (the heavy tails) occur more frequently than a normal distribution would predict, capturing actual market risk. This realism stems from our autonomous agents, which, based on BDI cognitive frameworks, process information heterogeneously, leading to diverse and potentially conflicting sentiments. These divergent sentiments then propagate through the integrated social network, becoming amplified and polarized, which in turn directly leads to extreme market outcomes.
Fat-tailed price changes: more extreme movements.
@@
Figure 4.2 shows TwinMarket capturing the leverage effect, where volatility responds asymmetrically. Specifically, initial price drops (x-axis Negative Returns) provoke larger subsequent drops (y-axis Lagged Negative Returns), thereby amplifying market volatility. This realistic asymmetry is a direct consequence of our LLM agents’ design; their BDI cognitive architectures are intentionally embedded with behavioral biases, such as loss aversion. This design makes them, akin to humans, inherently more sensitive to adverse signals. The integrated social media platform further amplifies the spread of negative sentiment, resulting in more widespread, synchronized, and potentially destabilizing collective trading behavior.
Initial price drops can trigger subsequent price drops.
@@
Figure 4.2 shows TwinMarket’s ability to replicate the volume-return relationship. Specifically, it shows that substantial increases in trading activity (x-axis Volume Change) are typically accompanied by significant market price fluctuations (y-axis Log Return). This co-movement is a key indicator, as in real markets, of collective behaviors such as herding among participants. In TwinMarket, the emergence of this co-movement provides compelling evidence demonstrates that our LLM agents, through their interactions, successfully generate collective behavioral patterns consistent with those observed in actual financial market dynamics.
Corr. between trading volume & price changes.
@@
Figure 4.2 shows TwinMarket successfully replicating volatility clustering: periods of large price swings (high Volatility on y-axis) tend to cluster together, as do periods of small swings (low Volatility). Both simulated and real volatility series clearly show these alternating groupings of high and low market turbulence over time. This emergent persistence arises as impactful market-moving events, processed via the BDI cognitive architecture, induce sustained heightened belief sensitivity. The social network then perpetuates these collective cognitive states by circulating related discussions or sentiments, prolonging periods of either erratic or stable trading.
Big price swings group together; small ones too.
@@
Real Data 7.26 0.14 p<0.01p{<}0.01 0.95
TwinMarket 5.24 0.11 p<0.01p{<}0.01 0.89
ABM-HPM 4.47 0.05 p<0.01p{<}0.01 0.82
ABM-BH 4.99 0.19 p<0.01p{<}0.01 0.72
We compare TwinMarket against two transitional agent-based models: the HPM Model [42], which captures investor strategy adaptation through herding behavior, predispositions, and reactions to price misalignments; and the BH Model [43, 21], which incorporates hyperbolic discounting and stochastic volatility to reflect heterogeneous investor behavior. As summarized in Table 4, TwinMarket consistently generates outputs that align more closely with real-world market statistics across all four stylized facts. This demonstrates its superior ability to replicate fundamental financial dynamics, thereby validating the empirical robustness of our simulation framework. These results were further validated for reproducibility and stability, as detailed in Appendix B.3.
The four stylized facts (SF) in Table 4 are quantified as follows: (1) Kurtosis (SF I): Measures tail heaviness; values greater than 3 indicate fat-tailed return distributions. (2) Negative Return Autocorrelation (SF II): Captures asymmetric responses to negative shocks. (3) Volume Return Correlation (SF III): A significant $p$-value (<0.01<0.01) suggests stronger volume–volatility linkage. (4) GARCH Volatility Clustering (SF IV): The sum $\alpha+\beta\approx 1$ reflects persistent volatility, a hallmark of financial time series [44].
@@ 5.2 Macro-level Emergence: Information Propagation
Previous work highlights how information propagation produces market‐wide echo effects [47, 48]. To probe the same mechanism, we identify the day’s most salient news items and deliver them directly to users with the highest network centrality, reproducing the information asymmetry characteristic of real markets. We then rerun the experiment with a parallel set of exaggerated, rumor-like headlines (We list some examples in Appendix E.4.) to stress-test the system. Observing how these targeted, micro-level interventions cascade through the network and reshape aggregate prices and volumes allows us to treat information diffusion as a source of macro-level emergence: the collective dynamics that arise cannot be inferred by examining any single agent in isolation, but only by tracking the system-wide response to the injected signals.
Figure 8: Heatmap of the behavioral similarity adjacency matrix. Higher density indicates greater behavioral similarity.
Intimation & Polarization. In TwinMarket, graph edges represent behavioral similarity, meaning users closely connected exhibit stronger intimation effects. As shown in Figure 7, high-degree users receive more upvotes, indicating their greater influence within the network. These opinion leaders [49] shape decision-making, fostering homogeneity and leading to polarization, as depicted in Figure 8 (further explored in Appendix B.4). The network’s adjacency matrix highlights clusters of trading similarity, further intensified under rumor conditions. This suggests that biased or extreme information spreads more rapidly in tightly connected groups, reinforcing polarization [50, 51].
@@ 6 Ablation Study
Table 5: Ablation study on the effects of the BDI framework and agent heterogeneity. $\uparrow$: higher is better, $\downarrow$: lower is better. Corr., RMSE, and MAE are computed on the daily normalized index price series between simulated outcomes and real-world data.
Model Variant RMSE $\downarrow$ MAE $\downarrow$ Corr. $\uparrow$ Kurtosis $\uparrow$ Lev. Effect $\uparrow$ GARCH $\alpha$ $\uparrow$ GARCH $\beta$ $\uparrow$
@@
The BDI ablation notably weakens the model’s ability to generate stylized facts, such as fat tails and asymmetric volatility, highlighting the importance of cognitive reasoning in agent behavior. In contrast, removing heterogeneity leads to a sharper collapse in realism, with near-Gaussian returns and negative correlation, underscoring the necessity of strategic diversity for emergent market complexity. Together, these results confirm that both BDI framework and agent heterogeneity are essential for replicating real-world market dynamics. Additional ablation studies appear in Appendix B.2.